OpenStudy (anonymous):

@Michele_Laino A jar contained 12 red marbles, 16 green marbles, and 10 blue marbles. A marble was selected at random, the color was recorded, and the marble was placed back in the jar. The table shows the results after 80 trials. What was the relative frequency of selecting a green marble? Round to the nearest percent, if necessary. Outcome Red Green Blue Number of trials 28 30 22

OpenStudy (michele_laino):

here the requested relative frequency f, is given by the subsequent ratio: \[f = \frac{{30}}{{28 + 30 + 22}} = ...?\]

OpenStudy (anonymous):

3/8

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

i have to round to the nearest percent

OpenStudy (michele_laino):

ok! than, you have to compute this: \[f = \frac{3}{8} \times 100 = ...?\]

OpenStudy (michele_laino):

oops...then*

OpenStudy (anonymous):

75/2

OpenStudy (michele_laino):

better is 38%, I think